3x/x-1 + 6x-9/x-1 -6 = 0









Answer by mangopeeler07(462)

Common denominator=x-1

3x/x-1 + 6x-9/x-1 -6 = 0
becomes

3x/x-1 + 6x-9/x-1 -6(x-1)/x-1=0

Distribute the -6

3x/x-1 + 6x-9/x-1 -6x+6/x-1=0

Combine the numerators
3x+6x-9-6x+6/x-1=0

Combine like terms in the numerator
3x-3/x-1=0

0 over anything is zero, so the numerator must be equal to zero. Eliminate the denominator.

3x-3=0
3x=3
x=1

Notice if you plug this back into the original equation, or any of the ones before you eliminated the denominator, the denominator becomes 0. That's undefined. So there is no solution.